40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 29 19:23:49 2013, Gabriele, Jenne, Frogs, LSC, Analog whitening filter of REFL55 not switching 
    Reply  Fri Mar 29 19:58:24 2013, Gabriele, Jenne, Frogs, LSC, Analog whitening filter of REFL55 not switching 
    Reply  Mon Apr 1 16:24:09 2013, Jamie, Frogs, LSC, PD whitening switching fixed (loose connection at break-out box) 20130401_161858.jpg20130401_161917.jpg
Message ID: 8383     Entry time: Mon Apr 1 16:24:09 2013     In reply to: 8375
Author: Jamie 
Type: Frogs 
Category: LSC 
Subject: PD whitening switching fixed (loose connection at break-out box) 

Quote:

We discovered that the analog whitening filter of the REFL55_I board is not switching when we operate the button on the user interface. We checked with the Stanford analyzer that the transfer function always correspond to the whitening on.

This turned out to just be a loose connection of the ribbon cable from Contec board in the LSC IO chassis at the BIO break-out box.  The DSUB connector at the break-out box was not strain relieved!  I reseated the connector and strain relieved it and now everything is switching fine.

20130401_161917.jpg

20130401_161858.jpg

I wonder if we'll ever learn to strain relieve...

ELOG V3.1.3-